### v4

@@ 3 Applications: Counting Defiers in Health Care
If a second stage outcome such as actual organ donation is available, under an assumption of no never takers, the instrumental variable estimate is no longer interpretable as an average treatment effect on compliers without further assumptions. As discussed by angrist1996, under an additional assumption that average treatment effects are the same for compliers and defiers, the instrumental variable estimate is again interpretable as a LATE. dechaisemartin2017 and tchetgen2024 propose alternative assumptions to interpret the instrumental variable estimand in the presence of compliers and defiers.
@@
The value of an assumption of no never takers becomes apparent if data are available on covariates. Under the assumption of no never takers, it is possible to label some specific people as compliers and others as defiers: all people in intervention who do not take up must be compliers, and all people who do not take up in control must be defiers. This reasoning allows us to make claims that the intervention had a causal effect on some specific people in the sample. Using terminology from pearl1999, under an assumption of no never takers, the intervention was a “necessary” cause of no takeup for all 11 people assigned intervention who do not take up, and was a “sufficient” cause of no takeup for all 31 people assigned control who do not take up. Under a monotonicity assumption, it is not possible to directly label any specific people as compliers or defiers because people who take up in intervention can be compliers or always takers, and people who do not take up in control can be compliers and never takers. An assumption that allows us to label some specific people as compliers and others as defiers could be an important first step to learn their characteristics, improve targeting in future samples, and obtain a larger average effect.
@@ 4 Contributions to the Literature
Our primary insight is that a randomized experiment can reveal evidence on an estimand more primitive than the average effect: the joint distribution of potential outcomes in the sample. The evidence can support a monotonicity assumption or inform a specific alternative assumption. We obtain evidence using the structure of the implemented randomization process to construct a design-based causal model and its resulting likelihood function, and we evaluate this evidence in a statistical decision theory framework. Our work contributes to the literature on design-based econometrics, statistical decision theory, and causal inference without monotonicity.
@@
In his 1935 book on the design of experiments, Fisher notes that the design of Charles Darwin’s experiments was “greatly superior” to the “statistical methods available at the time” (fisher1935). In their 2017 handbook chapter on the econometrics of experiments, Susan Athey and Guido Imbens write that they “recommend using statistical methods that are directly justified by randomization, in contrast to the more traditional sampling-based approach that is commonly used in econometrics” (athey2017). The sampling-based approach assumes that people in an experiment are randomly sampled from a population, and it attempts to learn about the population. The design-based approach, which we adopt, attempts to learn about the sample of people in the experiment, which is useful since participation in experiments and clinical trials can be far from random (alsan2025) and some samples do not come from a meaningful population (abadie2020).
@@
Our use of a design-based model rather than a sampling-based model enables our contributions. Our ability to learn about defiers with a design-based likelihood may seem paradoxical because it is well known that the traditional sampling-based likelihood does not vary with the number of defiers within the Fréchet bounds determined by the estimated marginal distributions of potential outcomes in intervention and control. As we review in Appendix B, for a given pair of marginal distributions, the likelihood function in the traditional sampling-based model is flat over all copulas connecting them. A large literature has focused on specifying what we can learn from the nonparametric Fréchet bounds in the traditional sampling-based model (balke1997; heckman1997; manski1997mixing; tian2000; zhang2003; imbens2004; fan2010; mullahy2018; li2019; bai2024; semenova2024).
The design-based model allows us to construct estimates that vary within the Fréchet bounds, as in Figure 1, by harnessing the parametric structure that follows directly from the randomization process. Parametric structure within randomized experiments is easy to justify as long as it corresponds to the actual randomization process, which is under the control of the people who design and implement the experiment. In many applications, economists use the parameteric structure of a model to infer aggregate counterfactuals from observed individual-level heterogeneity. We invert this procedure: using aggregate counterfactuals observed from the experiment, we apply the parametric structure of the randomization process to infer individual-level heterogeneity.
@@
Our use of the likelihood implied by the design-based model is central to our contributions. The design-based model defines counterfactuals using causal models of potential outcomes from neyman1923, welch1937, kempthorne1952, copas1973, rubin1974; rubin1977, greenland1986, holland1986, and others, and then uses the randomization structure to derive the design-based likelihood. While copas1973 derives the design-based likelihood and uses it to test hypotheses about the average effect and the Fisher null hypothesis, our advance is to use the likelihood to learn about a more primitive object: the joint distribution of potential outcomes in the sample, including the number of defiers. We revisit Copas’ design-based model in light of modern causal models, modern assumptions about the absence of defiers, and modern computational advances. Likelihood-based models can provide exact distributions of experimental data, test statistics, and confidence intervals, but they remain uncommon relative to simulation methods in randomization inference. One exception is ding2019, who use the design-based likelihood to examine the sensitivity of inference on aggregate parameters to assumptions about the joint distribution of potential outcomes. rigdon2015 and li2016 construct exact confidence intervals without the design-based likelihood.
@@
Our use of a decision-theoretic rule in the style of wald1949, rather than a classical hypothesis test, facilitates our progress. Classical hypothesis tests specify a null hypothesis and place the burden of evidence on the alternative hypothesis. This asymmetric treatment of Type I and Type II errors can imply an unrealistically conservative decision maker (tetenov2012) and exclude statistical evidence that, while weak, is still informative. To make use of all the evidence in our design-based likelihood, we focus on the statistical decision problem of choosing the correct joint distribution of potential outcomes in the sample, rather than on testing hypotheses about a prespecified distribution under prespecified size and power. We do not claim to provide a “test” of a monotonicity assumption. We provide a “decision” about the joint distribution of potential outcomes that can provide evidence for or against a monotonicity assumption.
@@
Our use of design-based statistical decision theory contributes to the integration of statistical decision theory into econometrics (manski2004; dehejia2005; manski2007; hirano2008; hirano2009; stoye2012; kitagawa2018; manski2018; manski2019; manski2021; fernandez2024). Econometrics is a natural discipline to embrace statistical decision theory because statistical decision theory has such a tight relationship with economic theory. Statistical decision theory is natural to apply in our design-based setting because it does not require large sample assumptions or asymptotic approximations. Our use of design-based statistical decision theory departs from “asymptotic analysis of statistical decision rules in econometrics” as reviewed in hirano2020 in favor of the literature that focuses on statistical decision theory in finite sample settings (canner1970; manski2007admissible; schlag2007; stoye2007; stoye2009; tetenov2012).
@@
We justify our proposed decision rule with justifications commonly employed in statistical decision theory, starting with the justification that our decision varies systematically with the data, as we show in Figure 3. One challenge with statistical decision criteria such as minimax and minimax regret is that researchers have shown instances in which they yield rules that do not vary with the data (schlag2003; manski2004; hirano2009; stoye2009). In his influential introduction to statistical decision theory, ferguson1967 recognizes that because maximum likelihood rules vary with the data, they are “reasonable” because they are better than “just guessing.” “However,” he continues, “decision theory, as developed here, is devoted to the problem of finding optimal rules, so that we do not refer to maximum likelihood estimates […] unless they turn out naturally to be optimal in some sense.”
@@
In Section 2.4.2, we further justify our decision rule as Bayes optimal under the appropriate conditions, which advances Bayesian decision theory (chamberlain2011) by incorporating the design-based likelihood for our class of experiments. Bayes optimality implies that our rule is admissible—that is, there is no alternative rule that performs at least as well for all possible joint distributions of potential outcomes, and strictly better for some (ferguson1967). Furthermore, it ensures that our rule cannot be bested in a betting framework (freedman1969). The Bayesian interpretation of our rule facilitates our auxiliary construction of credible sets on the joint distribution of potential outcomes in Table E.1, providing an exact Bayesian alternative to classical asymptotic inference on Fréchet bounds (manski1992; horowitz2000; tamer2004; jun2023) and the number of defiers within the Fréchet bounds (imbens2004).
@@
In addition to standard decision theoretic justifications, we also offer a perspective that our design-based maximum likelihood decision rule is optimal through the principle of maximum entropy (jaynes1957a; jaynes1957b), which does not require specification of a prior. golan2002 reviews a literature in econometrics that uses the principle of maximum entropy to abstract away from functional form assumptions on the likelihood function. In our setting, the design of the randomization, which is under the control of the experimenter, determines the functional form of the likelihood function, which also determines the functional form of the entropy.
@@
Our results provide evidence that can support a monotonicity assumption or a specific alternative, which can be useful because monotonicity assumptions have become widespread, but they are not always realistic. Influential work by imbens1994 proposes a monotonicity assumption that assumes away compliers or defiers in the first stage of an instrumental variable model, and manski1997monotone proposes an analogous assumption in the reduced form. Monotonicity assumptions are ubiquitous in the analysis of first stages, and they are gaining traction in the analysis of reduced forms (alsan2025). First stage monotonicity assumptions are useful for the interpretation of instrumental variable estimates as local average treatment effects on compliers, and we view them as reasonable in many contexts. However, they can be difficult to defend in some contexts, particularly in health care settings that motivate our work because interventions may be helpful for some but harmful for others. We interpret our binary outcome as “takeup” in our stylized example and applications so that we can use the familiar terminology of always and never “takers” developed for the first stage (angrist1996), but the design-based likelihood is valid for other binary outcomes, like survival. Given concerns about side effects in medicine, we see potential for useful applications that have perfect compliance and a single binary outcome that represents survival. In those applications, evidence of defiers could motivate researchers to reduce the dosage of a potentially toxic intervention.
@@
Previous approaches have revealed evidence that violates monotonicity assumptions, but only with additional data beyond our binary intervention and outcome. Machine learning approaches of wager2018 and semenova2024 require data on covariates to reveal heterogeneous effects. Analysis of side effects in medicine requires data on secondary outcomes (e.g. bernard2001). Specification tests of instrumental variable model assumptions can reveal evidence against monotonicity in the first stage (imbens1997; richardson2010; huber2012; huber2015testing; kitagawa2015; mourifie2017; machado2019), and marginal treatment effect methods can reveal evidence against monotonicity in the second stage (bjorklund1987; heckman1999; kowalski2023behavior; kowalski2023reconciling), but such approaches require the additional structure of a two-stage model as well as data on both stages. Evidence against first stage monotonicity by chan2022 requires a multi-valued instrument and the ability to observe a second stage outcome for one value of the first stage outcome. We contribute to the literature on violations of monotonicity by demonstrating that, even without additional data, our design-based model of an experiment can reveal evidence for or against monotonicity.
@@
We view the development of our visualizations as a secondary contribution. Figure 1 allows us to depict how the same data can arise from different distributions of potential outcomes within a Fréchet set. Our visualization of how the MLE varies with the data in Figure 3 allows us to depict all possible results from an experiment of a given size, demonstrating that we have not engaged in selective reporting. The visualization allows us to characterize patterns that would be difficult to infer from isolated empirical applications and demonstrates the value of statistical decision theory over testing.
@@ 5 Implications for Econometric and Applied Research
We see our demonstration that a design-based model can reveal evidence beyond the average effect with very minimal assumptions and data as a proof-of-concept that could pave the way for the production of stronger evidence that will achieve the ultimate goal of personalizing decisions. Even if the average effect is the decision-relevant parameter in some contexts, learning about heterogeneity that underlies the average effect is a first step toward targeting that can ultimately yield a larger average effect. In our second empirical application, under the maximum likelihood joint distribution of potential outcomes, which includes no never takers, we can deduce that the people who are untreated in intervention must be defiers, and the people who are untreated in control must be compliers. Under monotonicity, it is never possible to deduce that a given individual is a complier, and defiers are assumed away. Our hope is that the ability to look directly at some compliers and defiers under alternative assumptions motivated by our statistical decision rule will pave the way to characterize them, target interventions toward compliers and away from noncompliers, and improve the average effect of future interventions.
@@
Our work suggests several areas for further econometric research. Design-based likelihoods can be derived for more complex experimental settings, including those with two stages, multi-valued variables, and alternative randomization designs, such as matched pair, stratified, randomized permuted block, and designs that allow for rerandomization. A promising direction is applying design-based likelihoods to parameters traditionally studied under the lens of partial identification, such as the average treatment effect in experiments with imperfect compliance, missing data, attrition, spillovers, and measurement error. Another is using design-based likelihoods and entropy to motivate alternative parametric assumptions in auctions (jun2024) and other designed markets (abdulkadiroğlu2017). Approximation methods and advances in algorithmic efficiency could facilitate calculations in a wide array of samples. Optimal estimators under alternative priors and utility functions could be implemented with straightforward modifications. We are particularly interested in extending decision problems with “do no harm” objectives (mueller2023; benmichael2024; christy2024a; koch2025; gelman2025) to incorporate evidence from the design-based likelihood. Estimators could also be developed for alternative estimands, like potential outcome quantiles (cui2023; guggenberger2024). Finally, while previous results on optimal experimental design focus on estimating the average effect (bai2022), other designs, especially those that incorporate additional data, could reveal stronger evidence on the joint distribution of potential outcomes.
@@
Design-based methods have been around for a long time, and we have demonstrated that it is feasible to implement our design-based decision rule, but a practical impediment to widespread implementation is that applied researchers often do not report enough information about their experiments to replicate them with design-based methods (young2019). We use large language models to pull (openAI’s GPT-4o-mini) and categorize (openAI’s GPT-o3-mini) the randomization processes from 2,080 papers associated with randomized controlled trials from the Abdul Latif Jameel Poverty Action Lab (J-PAL). Only 61% have enough description for the large language model to confidently categorize the randomization method from the paper or associated AEA RCT Registry entry. However, of those, 78% use designs captured by design-based likelihoods that we consider. In the full set of papers, over 60% describe some mechanism through which there could be defiers. Another practical impediment to design-based methods is that instead of reporting absolute numbers, several researchers only report rates or regression results. Mandated reporting of experimental design and standard sample counts by J-PAL, medical journals, and other authorities could facilitate the application of our design-based decision rule, allowing researchers to learn more from their experiments.
@@
Appendix A  Derivation of the Design-Based Likelihood
The design-based likelihood of the joint distribution of potential outcomes in the sample ${\boldsymbol{\theta}}=(\theta_{11},\theta_{10},\theta_{01},\theta_{00})$ given the data $\boldsymbol{x}$} is the probability of this realization of the data given $\boldsymbol{\theta}$}. First, we define the joint distribution of potential outcomes in the intervention arm ${\boldsymbol{I}}\equiv\big(I_{11},I_{10},I_{01},I_{00}\big)$ as a random vector whose elements represent the numbers of always takers, compliers, defiers, and never takers randomized into intervention; we write its realizations as ${\boldsymbol{i}}=(i_{00},i_{10},i_{01},i_{00})$. These values are unobserved to the experimenter. However, the randomization process—the design of the experiment—implies a data generating process for the distribution of potential outcomes in the intervention arm $\boldsymbol{I}$} conditional on the distribution of potential outcomes in the sample $\boldsymbol{\theta}$}, which we can then use to derive the distribution of the data $\boldsymbol{X}$}.

### v5

@@ 3 Applications: Counting Defiers in Health Care
If a second stage outcome such as actual organ donation is available, under an assumption of no never takers, the instrumental variable estimate is no longer interpretable as an average treatment effect on compliers without further assumptions. As discussed by Angrist, Imbens, and Rubin (1996), under an additional assumption that average treatment effects are the same for compliers and defiers, the instrumental variable estimate is again interpretable as a LATE. De Chaisemartin (2017) and Tchetgen Tchetgen (2024) propose alternative assumptions to interpret the instrumental variable estimand in the presence of compliers and defiers.
@@
Although the design-based maximum likelihood estimates target the distribution of potential outcomes in the sample rather than the population, an evidence-based assumption of no never takers in the sample can facilitate learning about defiers in the population if data are available on covariates. Under the assumption of no never takers, it is possible to label some specific people in the sample as compliers and others as defiers: all people in intervention who do not take up must be defiers, and all people who do not take up in control must be compliers. This reasoning allows us to make claims that the intervention had a causal effect on some specific people in the sample. Using terminology from Pearl (1999), under an assumption of no never takers, the intervention was a “necessary” cause of no takeup for all 11 people assigned intervention who do not take up, and it would have been a “sufficient” cause of takeup for all 31 people assigned control who do not take up. Under a monotonicity assumption, it is not possible to directly label any specific people as compliers or defiers because people who take up in intervention can be compliers or always takers, and people who do not take up in control can be compliers or never takers. An assumption that allows us to label some specific people as compliers and others as defiers could be an important first step to learn their characteristics, improve targeting in future samples, and obtain a larger average effect.
@@ 4 Contributions to the Literature
Our primary insight is that a randomized experiment can reveal evidence on an estimand more primitive than the average effect: the joint distribution of potential outcomes in the sample. The evidence can support a monotonicity assumption or inform a specific alternative assumption. We obtain evidence using the structure of the implemented randomization process to construct a design-based causal model and its resulting likelihood function. Our work contributes to the literature on design-based econometrics, statistical decision theory, and causal inference without monotonicity.
@@
In his 1935 book on the design of experiments, Fisher notes that the design of Charles Darwin’s experiments was “greatly superior” to the “statistical methods available at the time” (Fisher, 1935). In their 2017 handbook chapter on the econometrics of experiments, Susan Athey and Guido Imbens write that they “recommend using statistical methods that are directly justified by randomization, in contrast to the more traditional sampling-based approach that is commonly used in econometrics” (Athey and Imbens, 2017). The sampling-based approach assumes that people in an experiment are randomly sampled from a population, and it attempts to learn about the population. The design-based approach, which we adopt, attempts to learn about the sample of people in the experiment, which is useful since participation in experiments and clinical trials can be far from random (Alsan, Durvasula, Gupta, Schwartzstein, and Williams, 2024) and some samples do not come from a meaningful population (Abadie, Athey, Imbens, and Wooldridge, 2020).
@@
Our use of a design-based model rather than a sampling-based model enables our contributions. Our ability to learn about defiers with a design-based likelihood may seem paradoxical because it is well known that the traditional sampling-based likelihood does not vary with the number of defiers within the Fréchet bounds determined by the estimated marginal distributions of potential outcomes in intervention and control, as we review in Appendix A.2. Furthermore, it is straightforward to show that the Fréchet bounds on defiers always include zero if the estimated average effect shows compliers on average. A large literature has focused on specifying what we can learn from the nonparametric Fréchet bounds in the traditional sampling-based model (Balke and Pearl 1997; Heckman, Smith, and Clements 1997; Manski 1997b; Tian and Pearl 2000; Zhang and Rubin 2003; Imbens and Manski 2004; Fan and Park 2010; Mullahy 2018; Li and Pearl 2019; Bai, Huang, Moon, Shaikh, and Vytlacil 2024; Semenova 2024).
@@
Our use of the likelihood implied by the design-based model is central to our contributions. The design-based model defines counterfactuals using causal models of potential outcomes from Neyman (1923), Welch (1937), Kempthorne (1952), Copas (1973), Rubin (1974, 1977), Greenland and Robins (1986), Holland (1986), and others, and then uses the randomization structure to derive the design-based likelihood. While Copas (1973) derives the design-based likelihood and uses it to test hypotheses about the average effect and the Fisher null hypothesis, our advance is to use the likelihood to learn about a more primitive object: the joint distribution of potential outcomes in the sample, including the number of defiers. We revisit Copas’ design-based model in light of modern causal models, modern assumptions about the absence of defiers, and modern computational advances. Likelihood-based models can provide exact distributions of experimental data, test statistics, and confidence intervals, but they remain uncommon relative to simulation methods in randomization inference. One exception is Ding and Miratrix (2019), who use the design-based likelihood to examine the sensitivity of inference on aggregate parameters to assumptions about the joint distribution of potential outcomes. Rigdon and Hudgens (2015) and Li and Ding (2016) construct exact confidence intervals without the design-based likelihood. We use the design-based likelihood for estimation as well as inference.
While the design-based model provides evidence about defiers beyond the Fréchet bounds in the sample, it is important to consider what evidence, if any, it can provide about the population from which the sample was drawn. Sometimes, evidence about the population could be more relevant for policy than evidence about the sample. Under standard sampling assumptions, the experiment does not provide evidence about defiers beyond the Fréchet bounds in the population. While standard sampling models assume that the data consist of independent and identically distributed random draws from a population, this assumption is often far from accurate—especially for small, finite populations sampled without replacement. When the population is finite and the sample is drawn without replacement, departure from standard sampling assumptions can propagate information about defiers in the sample to the population. Suppose that the 6 people in our stylized example were randomly drawn without replacement from a finite population of 24 people to participate in a pilot study. As we show in Appendix A.3, it is possible to derive an alternative sampling-based likelihood of the joint distribution of potential outcomes in the finite population under sampling without replacement, and it can reveal evidence within the Fréchet bounds. In our stylized example, the maximum likelihood estimate for the finite population of 24 based on an experimental sample that was one fourth of the population includes 16 compliers and 8 defiers, exactly four times the estimate in the sample. In terms of likelihood ratios, the evidence revealed within the finite sample is weak, and the evidence revealed within the finite population is even weaker. Furthermore, as the finite population grows, the evidence gets weaker until the finite population becomes infinite, and there is no evidence within the Fréchet bounds. Nevertheless, weak evidence can improve decision making over no evidence at all.
@@
Our engagement with statistical decision theory in the style of Wald (1949), rather than classical hypothesis tests, facilitates our progress. Classical hypothesis tests specify a null hypothesis and place the burden of evidence on the alternative hypothesis. This asymmetric treatment of Type I and Type II errors can imply an unrealistically conservative decision maker (Tetenov, 2012) and exclude statistical evidence that, while weak, is still informative. To make use of weak evidence from our design-based likelihood, we focus on the statistical decision problem of choosing the correct joint distribution of potential outcomes in the sample, rather than on testing hypotheses about a prespecified distribution under prespecified size and power. We do not claim to provide a “test” of a monotonicity assumption. We provide a “decision” about the joint distribution of potential outcomes that can provide evidence for or against a monotonicity assumption.
@@
Our use of design-based statistical decision theory contributes to the integration of statistical decision theory into econometrics (Manski, 2004; Dehejia, 2005; Manski, 2007; Hirano, 2008; Hirano and Porter, 2009; Stoye, 2012; Kitagawa and Tetenov, 2018; Manski, 2018, 2019; Manski and Tetenov, 2021; Fernández, Montiel Olea, Qiu, Stoye, and Tinda, 2024). Econometrics is a natural discipline to embrace statistical decision theory because statistical decision theory has such a tight relationship with economic theory. Statistical decision theory is natural to apply in our design-based setting because it does not require large sample assumptions or asymptotic approximations. Our use of design-based statistical decision theory departs from “asymptotic analysis of statistical decision rules in econometrics” as reviewed in Hirano and Porter (2020) in favor of the literature that focuses on statistical decision theory in finite sample settings (Canner, 1970; Manski and Tetenov, 2007; Schlag, 2007; Stoye, 2007, 2009; Tetenov, 2012).
@@
We justify our maximum likelihood estimates with justifications commonly employed in statistical decision theory, starting with the justification that our estimates vary systematically with the data, as we show in Figure B.1. One challenge with statistical decision criteria such as minimax and minimax regret is that researchers have shown instances in which they yield rules that do not vary with the data (Schlag, 2003; Manski, 2004; Hirano and Porter, 2009; Stoye, 2009). In his influential introduction to statistical decision theory, Ferguson (1967) recognizes that because maximum likelihood rules vary with the data, they are “reasonable” because they are better than “just guessing.” “However,” he continues, “decision theory, as developed here, is devoted to the problem of finding optimal rules, so that we do not refer to maximum likelihood estimates […] unless they turn out naturally to be optimal in some sense.”
@@
In Appendix B.3, we further justify our maximum likelihood estimates as Bayes optimal under the appropriate conditions, which advances Bayesian decision theory (Chamberlain, 2011) by incorporating the design-based likelihood. Bayes optimality implies that our rule is admissible—that is, there is no alternative rule that performs at least as well for all possible joint distributions of potential outcomes, and strictly better for some (Ferguson, 1967). Furthermore, it ensures that our rule cannot be bested in a betting framework (Freedman and Purves, 1969). The Bayesian interpretation of our estimates facilitates our auxiliary construction of credible sets on the joint distribution of potential outcomes in Table C.3, providing an exact Bayesian alternative to classical asymptotic inference on Fréchet bounds (Manski, Sandefur, McLanahan, and Powers, 1992; Horowitz and Manski, 2000; Tamer, Chernozhukov, and Hong, 2004; Jun and Lee, 2023) and the number of defiers within the Fréchet bounds (Imbens and Manski, 2004).
@@
In addition to standard decision theoretic justifications, we also offer a perspective that our design-based maximum likelihood estimates are optimal through the principle of maximum entropy (Jaynes, 1957b, a), which does not require specification of a prior. Golan (2002) reviews a literature in econometrics that uses the principle of maximum entropy to abstract away from functional form assumptions on the likelihood function. In our setting, the design of the randomization, which is under the control of the experimenter, determines the functional form of the likelihood function, which also determines the functional form of the entropy.
@@
Our results provide evidence that can support a monotonicity assumption or a specific alternative, which can be useful because monotonicity assumptions have become widespread, but they are not always realistic. Influential work by Imbens and Angrist (1994) proposes a monotonicity assumption that assumes away compliers or defiers in the first stage of an instrumental variable model, and Manski (1997a) proposes an analogous assumption on treatment response. Monotonicity assumptions are ubiquitous in the analysis of first stages, and they are gaining traction elsewhere (Alsan, Cawley, Doyle, and Skelley, 2025). First stage monotonicity assumptions are useful for the interpretation of instrumental variable estimates as local average treatment effects on compliers, and we view them as reasonable in many contexts. However, they can be difficult to defend in some contexts, particularly in health care settings that motivate our work because interventions may be helpful for some but harmful for others. We interpret our binary outcome as “takeup” in our stylized example and applications so that we can use the familiar terminology of always and never “takers” developed for the first stage (Angrist, Imbens, and Rubin, 1996), but the design-based likelihood is valid for other binary outcomes, like survival. Given concerns about side effects in medicine, we see potential for useful applications that have a single binary outcome that represents survival. In those applications, evidence of defiers could motivate researchers to reduce the dosage of a potentially toxic intervention.
@@
Previous approaches have revealed evidence that violates monotonicity assumptions, but only with additional data beyond our binary intervention and outcome. Machine learning approaches of Wager and Athey (2018) and Semenova (2024) require data on covariates to reveal heterogeneous effects. Analysis of side effects in medicine requires data on secondary outcomes (e.g. Bernard et al., 2001). Specification tests of instrumental variable model assumptions can reveal evidence against monotonicity in the first stage (Imbens and Rubin, 1997; Richardson and Robins, 2010; Huber and Mellace, 2012, 2015; Kitagawa, 2015; Mourifié and Wan, 2017; Machado, Shaikh, and Vytlacil, 2019), and marginal treatment effect methods can reveal evidence against monotonicity in the second stage (Björklund and Moffitt, 1987; Heckman and Vytlacil, 1999; Kowalski, 2023b, a), but such approaches require the additional structure of a two-stage model as well as data on both stages. Evidence against first stage monotonicity by Chan, Gentzkow, and Yu (2022) requires a multi-valued instrument and the ability to observe a second stage outcome for one value of the first stage outcome. We contribute to the literature on violations of monotonicity by demonstrating that, even without additional data, our design-based model of an experiment can reveal evidence for or against monotonicity.
@@
We view the development of our visualizations as a secondary contribution. Figure 1 allows us to depict how the same data can arise from different distributions of potential outcomes within a Fréchet set. Our visualization of how the MLE varies with the data in Figure B.1 allows us to depict all possible results from an experiment of a given size, demonstrating that we have not engaged in selective reporting. The visualization allows us to characterize patterns that would be difficult to infer from isolated empirical applications and demonstrates the value of statistical decision theory over testing.
@@ 5 Implications for Econometric and Applied Research
We see our demonstration that a design-based model can reveal evidence beyond the average effect with very minimal assumptions and data as a proof-of-concept that could pave the way for the production of stronger evidence that will achieve the ultimate goal of personalizing decisions. Even if the average effect is the decision-relevant parameter in some contexts, learning about heterogeneity that underlies the average effect is a first step toward targeting that can ultimately yield a larger average effect. In our second empirical application, under the maximum likelihood joint distribution of potential outcomes, which includes no never takers, we can deduce that the people who are untreated in intervention must be defiers, and the people who are untreated in control must be compliers. Under monotonicity, it is never possible to deduce that a given individual is a complier, and defiers are assumed away. Our hope is that the ability to look directly at some compliers and defiers under alternative assumptions motivated by our estimates will pave the way to characterize them, target interventions toward compliers and away from noncompliers, and improve the average effect of future interventions.
Our findings have already inspired new research from statisticians and econometricians. In a previous version of our paper (Christy and Kowalski, 2024a), we engage directly with utility functions that depend on the joint distribution of potential outcomes, motivated by our work on mammograms, which can help some women but harm others through overdiagnosis (Kowalski, 2023b, a). Our work contributes to the literature on “asymmetric counterfactual utilities” (Li and Pearl, 2019; Mueller and Pearl, 2023; Ben-Michael, Imai, and Jiang, 2024). Alongside the literature, it has prompted engagement from Koch and Imai (2025) and inspired Gelman and Mikhaeil (2025). Our work has also inspired Chen, Roth, and Spiess (2026), which shows that the joint distribution of potential outcomes is formally identified using the design-based likelihood. Despite identification, like us, they emphasize that the potential for learning is limited. They find that there exist Bayesian priors that do not update about monotonicity under any realization of the data. In Appendix D, we demonstrate that while such priors exist, they are rare, and those they construct assign zero prior weight to many distributions, violating Cromwell’s rule (Lindley, 1985). A challenge for future research is to explore how best to use weak evidence on defiers in the sample and how to make it stronger.
@@
Our work suggests several areas for further methodological research. Design-based likelihoods can be derived for more complex experimental settings, including those with two stages, multi-valued variables, and alternative randomization designs, such as matched pair, stratified, randomized permuted block, and designs that allow for rerandomization. A promising direction is applying design-based likelihoods to parameters traditionally studied under the lens of partial identification, such as the average treatment effect in experiments with imperfect compliance, missing data, attrition, spillovers, and measurement error. Another is using design-based likelihoods and entropy to motivate alternative parametric assumptions in auctions (Jun and Pinkse, 2024) and other designed markets (abdulkadiroğlu2017). Approximation methods and advances in algorithmic efficiency could facilitate calculations in a wide array of samples. Optimal estimators under alternative priors and utility functions could be implemented with straightforward modifications. Estimators could also be developed for alternative estimands, like asymmetric counterfactual utilities, measures of discrimination from audit studies (Kline and Walters, 2021), and potential outcome quantiles (Cui and Han, 2023; Guggenberger, Mehta, and Pavlov, 2024). Finally, while previous results on optimal experimental design focus on estimating the average effect (Bai, 2022), other designs, especially those that incorporate additional data, could reveal stronger evidence on the joint distribution of potential outcomes.
@@
Design-based methods have been around for a long time, and we have demonstrated that it is feasible to implement our design-based decision rule, but a practical impediment to widespread implementation is that applied researchers often do not report enough information about their experiments to replicate them with design-based methods (Young, 2019). We use large language models to pull (OpenAI’s GPT-4o-mini) and categorize (OpenAI’s GPT-o3-mini) the randomization processes from 2,080 papers associated with randomized controlled trials from the Abdul Latif Jameel Poverty Action Lab (J-PAL). Only 61% have enough description for the large language model to confidently categorize the randomization method from the paper or associated AEA RCT Registry entry. However, of those, 78% use designs captured by design-based likelihoods that we consider. In the full set of papers, over 60% describe some mechanism through which there could be defiers. Another practical impediment to design-based methods is that instead of reporting absolute numbers, several researchers only report rates or regression results. Mandatory reporting of experimental design and standard sample counts by J-PAL, medical journals, and other authorities could facilitate the application of our design-based likelihood, allowing researchers to learn more from their experiments.
@@
Appendix A  Derivations of Likelihoods
Appendix A.1  Design-Based Likelihoods Can Vary Conditional on the Marginal Distributions of Potential Outcomes
The design of the experiment—the randomization process—allows us to derive the probability of a realization of the experimental data $\boldsymbol{x}$} given the joint distribution of potential outcomes in the sample $\boldsymbol{\theta}$}, which is equal to the likelihood $\mathcal{L}({\boldsymbol{\theta}}\mid{\boldsymbol{x}})$ by definition. Recall that the distribution of potential outcomes in the sample $\boldsymbol{\theta}$} consists of the number of always takers $\theta_{11}$, the number of compliers $\theta_{10}$, the number of defiers $\theta_{01}$, and the number of never takers $\theta_{00}$. We define the joint distribution of potential outcomes in the intervention arm ${\boldsymbol{I}}\equiv\big(I_{11},I_{10},I_{01},I_{00}\big)$ as a random vector whose elements represent the numbers of always takers, compliers, defiers, and never takers randomized into intervention; we write its realizations as ${\boldsymbol{i}}=(i_{11},i_{10},i_{01},i_{00})$. Using the randomization design, we can derive the distribution of these unobserved values and use it to compute the distribution of the data $\boldsymbol{X}$}.